OpenStudy (anonymous):

What is the distance (x) between any two holes on the rack? Each hole is 3/4 inches in diameter. the space at each end of rack is the same size as the space between any two holes.(there are 5 holes) and the rack is (12inches 3/4 ) long . if i missed any information Im sorry and please tell me what i missed.

OpenStudy (gw2011):

I assume the rack is 12 and 3/4 inches long. Since there are 5 holes and each hole is 3/4 inches, you get (5)(3/4)=15/4 inches=3 and 3/4 inches. Subtracting 3 and 3/4 inches from 12 and 3/4 inches, you are left with 9 inches. There are 6 spaces involved (4 spaces between the holes and a space at each end). Therefore, you divide the 9 inches left by the 6 spaces and you get 9/6=3/2=1 and 1/2 inches between each hole.

OpenStudy (anonymous):

:DDD thank you so much. i got the same answer too. i just wanted to make sure.i knew i wasn't loosing it

OpenStudy (anonymous):